# GreenIQ: A Deep Search Platform for Comprehensive Carbon Market Analysis and Automated Report Generation

### v1

@@ Abstract
This study introduces GreenIQ, an AI-powered deep search platform designed to revolutionise carbon market intelligence through autonomous analysis and automated report generation. Carbon markets operate across diverse regulatory landscapes, generating vast amounts of heterogeneous data from policy documents, industry reports, academic literature, and real-time trading platforms. Traditional research approaches remain labour-intensive, slow, and difficult to scale. GreenIQ addresses these limitations through a multi-agent architecture powered by Large Language Models (LLMs), integrating five specialised AI agents: a Main Researcher Agent for intelligent information retrieval, a Report Writing Agent for structured synthesis, a Final Reviewer Agent for accuracy verification, a Data Visualisation Agent for enhanced interpretability, and a Translator Agent for multilingual adaptation. The system achieves seamless integration of structured and unstructured information with AI-driven citation verification, ensuring high transparency and reliability. GreenIQ delivers a 99.2% reduction in processing time and a 99.7% cost reduction compared to traditional research methodologies. A novel AI persona-based evaluation framework involving 16 domain-specific AI personas highlights its superior cross-jurisdictional analytical capabilities and regulatory insight generation. GreenIQ sets new standards in AI-driven research synthesis, policy analysis, and sustainability finance by streamlining carbon market research. It offers an efficient and scalable framework for environmental and financial intelligence, enabling more accurate, timely, and cost-effective decision-making in complex regulatory landscapes
@@
  NOTE: Preprint, Accepted to ICMLT 2025, Helsinki, Finland
Bisola Faith Kayode is an Independent Researcher in the United Kingdom (e-mail: bisolakayode11@gmail.com). Akinyemi Sadeeq Akintola is with Universidade NOVA de Lisboa, Lisbon, Portugal (e-mail: sadeeq2@gmail.com). Oluwole Fagbohun is with Carbonnote, USA (e-mail: wole@readrly.io). Egonna Anaesiuba-Bristol is with the University of Surrey, United Kingdom (e-mail: egonnabristol@gmail.com). Onyekachukwu Ojumah is an Independent Researcher in the United Kingdom (e-mail: onyekaojumah22@gmail.com). Oluwagbade Odimayo is with the University of Essex, United Kingdom (e-mail: oluwagbadeodimayo@yahoo.com). Toyese Oloyede is an Independent Researcher in the United Kingdom (e-mail: toyesej@gmail.com). Aniema Inyang is with Vuhosi Limited, United Kingdom (e-mail: annywillow@gmail.com). Teslim Kazeem is an Independent Researcher in the United Kingdom (e-mail: teslimolukazeem@gmail.com). Habeeb Alli is with Readrly Limited, United Kingdom (e-mail: allihabeeboluwasegun@gmail.com). Udodirim Ibem Offia is with Vuhosi Limited, United Kingdom (e-mail: udodirim.offia@gmail.com). Prisca Chinazor Amajuoyi is with Vuhosi Limited, United Kingdom (e-mail: amajuoyiprisca1@gmail.com).

### v2

@@ Abstract
This study introduces GreenIQ, an AI-powered deep search platform designed to revolutionise carbon market intelligence through autonomous analysis and automated report generation. Carbon markets operate across diverse regulatory landscapes, generating vast amounts of heterogeneous data from policy documents, industry reports, academic literature, and real-time trading platforms. Traditional research approaches remain labour-intensive, slow, and difficult to scale. GreenIQ addresses these limitations through a multi-agent architecture powered by Large Language Models (LLMs), integrating five specialised AI agents: a Main Researcher Agent for intelligent information retrieval, a Report Writing Agent for structured synthesis, a Final Reviewer Agent for accuracy verification, a Data Visualisation Agent for enhanced interpretability, and a Translator Agent for multilingual adaptation. The system achieves seamless integration of structured and unstructured information with AI-driven citation verification, ensuring high transparency and reliability. GreenIQ delivers a 99.2% reduction in processing time and a 99.7% cost reduction compared to traditional research methodologies. A novel AI persona-based evaluation framework involving 16 domain-specific AI personas highlights its superior cross-jurisdictional analytical capabilities and regulatory insight generation. GreenIQ sets new standards in AI-driven research synthesis, policy analysis, and sustainability finance by streamlining carbon market research. It offers an efficient and scalable framework for environmental and financial intelligence, enabling more accurate, timely, and cost-effective decision-making in complex regulatory landscapes.
@@
  NOTE: Preprint, Submitted to ICMLT 2025, Helsinki, Finland
Oluwole Fagbohun is with Carbonnote, Maryland, USA (e-mail: wole@carbonnote.ai). Sai Yashwanth is with Vuhosi Limited, London, United Kingdom (e-mail: taddishetty34@gmail.com). Akinyemi Sadeeq Akintola is with Universidade NOVA de Lisboa, Lisbon, Portugal (e-mail: sadeeq2@gmail.com). Ifeoluwa Wurola is with the University of Hull, Hull, United Kingdom (e-mail: wuraolaifeoluwa@gmail.com). Lanre Shittu is with Vuhosi Limited, London, United Kingdom (e-mail: hr.lanreshittu@gmail.com). Aniema Inyang is with Readrly Limited, London, United Kingdom (e-mail: annywillow@gmail.com). Oluwatimilehin Odubola is with ResilStudio Ltd, London, United Kingdom (e-mail: timi@resilstudio.com). Udodirim Offia is with Readrly Limited, London, United Kingdom (e-mail: udodirim.offia@gmail.com). Said Olanrewaju is with Cnergy Capital Limited, London, United Kingdom (e-mail: said.olar@cnergyfund.com). Ogidan Toluwaleke is with Vuhosi Limited, London, United Kingdom (e-mail: gbemilekeogidan@gmail.com). Ilemona Abutu is with Carbonnote, Maryland, USA (e-mail: abutu.mona@gmail.com). Taiwo Akinbolaji is with Silveredge Consults, London, United Kingdom (e-mail: taiwoakinbolaji@gmail.com).